### v2

# Distributed acoustic sensing for ocean applications

Abstract
Extensive monitoring of acoustic activities is important for many fields, including biology, security, oceanography, and Earth science. Distributed acoustic sensing (DAS) is an evolving technique for continuous, wide-coverage measurements of mechanical vibrations, which is suited to ocean applications. DAS illuminates an optical fiber with laser pulses and measures the backscattered wave due to small random variations in the refractive index of the material. Specifically, DAS uses coherent optical interferometry to measure the phase difference of the backscattered wave from adjacent locations along the fiber. External stimuli, such as mechanical strain due to acoustic wavefields impinging on the fiber-optic cable, modulate the backscattered wave. Hence, the differential phase measurements of the optical backscatter are proportional to the underlying physical quantities of the surrounding wavefield. Continuous measurement of the backscattered electromagnetic signal provides a distributed sensing modality that extends spatially along the fiber. We provide an overview of DAS technology and detail the underlying physics, from electromagnetic to mechanical and eventually acoustic quantities. We explain the effect of DAS acquisition parameters in signal processing and show the potential of DAS for sound source detection on data collected from the Ocean Observatories Initiative Regional Cabled Array https://doi.org/10.58046/5J60-FJ89.
I Introduction
Light propagating in an optical fiber undergoes backscattering due to small random variations in the refractive index of the material. Fiber-optic sensing technology is based on the phase modulation of the backscattered light traveling in an optical fiber due to external stimuli such as temperature gradients or mechanical strain. Distributed sensing transforms a fiber cable into a very long, dense sensor arraygrattan2000fiber that can extend over hundredswaagaard2021realtime or even thousands of kilometers,mazur2024real ; ronnekleiv2025range offering unprecedented sensing coverage with rapidly advancing technology. Moreover, fiber-optic sensing makes temperature and vibration measurements in harsh environments accessible due to its robustness to electromagnetic interference, high pressure, and temperature.Miah2017 ; Lindsey2021
Distributed acoustic sensing (DAS) is a fiber-optic sensing technology, which illuminates an optical fiber with laser pulses and measures the phase of the backscattered wave along the fiber.Hartog2017 A fiber-optic cable that is in mechanical contact with the surrounding medium will be stressed by external mechanical vibrations. Strain, i.e., the relative deformation of the fiber under stress, alters the distance traveled by the backscattered light changing its phase.KennettBook DAS records the phase of the backscattered light to estimate the strain (or strain rate) along the fiber and infer the displacement (or velocity) of the mechanical vibration.masoudi2016contributed ; yang2022sub Elongated optical fibers are mostly sensitive to the axial component of the induced strain.martin2021introduction ; williams2024toward Consequently, without a tangled cable layout,kuvshinov2016interaction ; ning2018multicomponent DAS exhibits directional sensitivity to the measured mechanical wavefield.fang2023directional Given that stress and strain are related through the elastic modulus of the solid material according to Hooke’s law,Brekhovskikh1994 both the fiberglass coatingbudiansky1979pressure and the deployment conditions of the cable, e.g., burial depth and seafloor type,sladen2019distributed affect the strain sensitivity of the fiber optic cable to the applied stress.
DAS has gained momentum in geophysics as a cost-effective, high-resolution method for seismic monitoring and seismic tomography.mateeva2014distributed ; zhan2020distributed ; Lindsey2021 Applications range from near-surface seismic monitoring of ambient noisedou2017distributed ; saengduean2024multi to borehole measurements of hydraulic fracturingjin2017hydraulic ; chien2023automatic and monitoring the sea state dynamics with submarine cables.lindsey2019illuminating ; lin2024measurement ; chien2025calibration Recently, the potential of DAS technology has been leveraged for monitoring the underwater acoustic wavefield with promising results in the identification and localization of low-frequency sound sources such as blue and fin whale vocalizationsbouffaut2022eavesdropping ; landro2022sensing ; wilcock2023distributed and ship noise.rivet2021preliminary ; shao2025tracking ; paap2025leveraging Monitoring the ocean soundscape at low frequencieswilcock2014sounds is key to understanding the impact of anthropogenic noiseandrew2011long ; chapman2011low ; mcdonald2006increases and climate changegavrilov2008long on ecology,gavrilov2012steady security, and marine exploration.miksis2016low So far, the characterization of the ocean soundscape has been based on recordings from hydrophones and hydrophone arrays with limited spatial coverage.mcdonald2006increases ; chapman2011low DAS converts the length of the fiber cable to an extensive linear aperture, which is especially useful for localization of low-frequency sound sources (below 50-100 Hz).
Despite the increasing interest in DAS technology for high-resolution observations of acoustic signals, the current applications in ocean acoustics are limited to the aforementioned low-frequency wavefields, i.e., whale calls, ship noise, and ocean swell. The absence of ocean acoustic observations at higher frequencies is related to the acquisition parameters of the current DAS systems, which are set with the objective of a sufficient signal-to-noise ratio. Nevertheless, calibration experiments with controlled transmitted signalsdouglass2023distributed ; shen2024high and at-sea experiments for underwater communicationspotter2024first indicate that DAS detection capability is not limited to low-frequency sounds. This paper links the acquisition parameters with signal processing and provides insight for improved observations of acoustic signals.
After trend removal and band-pass filtering to the frequency range of interest, the data are commonly analyzed with interferometric methods,saengduean2024multi $f$-$k$ decomposition, and delay-and-sum beamforming.nasholm2022array Given the high dimensionality of DAS data, machine learning provides compelling methods for dimensionality reduction and feature clustering.chien2023automatic
In this paper, we provide an overview of DAS technology. We detail the mathematical derivation of the underlying physics, from electromagnetic to mechanical and eventually acoustic quantities in Section II. The mathematical analysis highlights the physical principles of DAS in a generic manner, aiming to inform the choice of data acquisition parameters rather than to detail the technological specificity of different systems, which are evolving rapidly. Section III conforms the spatial sensitivity of DAS to that of an array of spatially distributed sensors of a sizeable aperture. Section IV analyzes publicly available data for sound source detection based on the phase of the acoustic particle displacement. Finally, Section VII concludes the paper.
Table 1: List of parameters and operators.
Space Time Direction Operators
$z$ axial distance $t$ time $\sin\theta$ directional of impinging wave $\widetilde{[\cdot]}$ complex quantity
$L_{f}$ interrogated fiber length $\tau_{\text{max}}$ two-way travel time along fiber length $\theta$ polar angle $[\cdot]^{*}$ complex conjugation
$L_{p}$ light pulse length $T_{p}$ light pulse duration (width) $k_{z}$ wavenumber of acoustic wave along the axial dimension $\widehat{[\cdot]}$ estimated quantity
$L_{g}$ gauge length $\Delta\Phi$ differential phase of backscattered light $[\cdot]_{z}$ projection to the fiber’s axis
$L_{h}$ channel distance - distance between spatial samples $\tau_{\text{max}}$ distance between temporal samples $[\cdot]_{a}$ quantity related to the acoustic field
$f_{s}$ sampling frequency (spatial sampling) $f_{r}$ pulse repetition frequency (temporal sampling) [][]delimited-[]delimited-[][$\cdot$]\!*\![$\cdot$][ ⋅ ] ∗ [ ⋅ ] convolution
Figure 1: (color online) Schematic representation of scattering mechanisms due to a light pulse of frequency $f_{p}$ and wavelength $\lambda_{p}$ propagating through an optical fiber. In single-mode fibers, $f_{p}$ is typically about 200 THz ($\lambda_{p}\approx$ 1550 nm). External stimuli, such as an increase in temperature $T$ or strain $\epsilon$, affect the respective scattering response as indicated by the arrows. The horizontal axis represents the wavelength $\lambda$ or, in reverse direction the frequency $f$ of the backscattered light. Stokes processes refer to an up-shift of the wavelength of the scattered light $\lambda>\lambda_{p}$, whereas anti-Stokes to a down-shift $\lambda<\lambda_{p}$. The Stokes radiation for Raman scattering is relatively insensitive to temperature and has a lower intensity than the corresponding anti-Stokes radiation due to lower scattering efficiency at lower frequencies.Ukil2011
II Distributed Acoustic Sensing Technology
DAS measures the effect of external mechanical vibrations on the backscattering of laser light traveling along an optical fiber. This section details the DAS technology by (II.1) providing an overview of the light scattering mechanisms within a fiber optic cable, (II.2-II.3) deriving the mathematical expression for the backscattered wave under monochromatic excitation, (II.4) explaining the detection method of the backscattered light, (II.5) showing how the measurand is related to strain from external mechanical stress, and finally (II.6) explaining the typical acquisition parameters and the data structure. Table 1 provides a list of the important parameters for the analysis.
II.1 Scattering mechanisms
An optical fiber is a thin thread of silica glass with a core of dielectric material, clad with a lower refractive index material. The difference in refractive index between the core and the cladding creates an optical waveguide for light transmitted in the fiber. Optical fibers are designed to have a core with a constant refractive index such that light transmitted into the fiber travels undisturbed along the fiber. However, in practice, small density variations in the core material due to impurities or the manufacturing procedure introduce small random variations in the refractive index along the optical fiber, which cause scattering of the propagating electromagnetic wave.KennettBook Hence, the refractive index along the fiber’s axis can be described as
$n(z)=n_{f}+\Delta n(z),$ (1)
where $n_{f}$ is the nominal refractive index of the optical fiber’s core material and $\Delta n(z)$ describes the localized fluctuation of the refractive index. The refractive index determines the speed of light in the optical fiber as $c_{n}\approx c/n_{f}$, where $c\approx 3\times 10^{8}$ m/s is the speed of light in vacuum.
The main scattering mechanisms are Rayleigh, Brillouin, and Raman scattering.masoudi2016contributed Rayleigh scattering is an elastic process caused by local scatterers with dimensions much smaller than the wavelength that does not involve energy transfer to the scatterer and does not shift the frequency of the incident wave. Brillouin and Raman scattering are inelastic processes caused by changes in the vibrational states in a lattice and molecular level, respectively, that involve energy transfer to the scatterer and, as a result, shift the frequency of the incident wave.Cotter1983 ; Shiota1992 Brillouin scattering can be leveraged for both strain and temperature measurements,Bao2011 whereas Raman scattering can be leveraged for distributed temperature sensing.Ukil2011 Rayleigh scattering exhibits the highest intensity and is principally exploited for DAS measurements.KennettBook Figure 1 depicts the effect of external stimuli such as temperature and strain to elastic and inelastic optical scattering mechanisms [Grattan2000, , Ch. 3.8].
Figure 2: (color online) Schematic representation of active optic sensing for passive monitoring of external vibrational fields. (a) A single-mode fiber of total length $L_{f}$ and a small cross-section constructed with randomly distributed impurities in the core material. (b) The fiber is interrogated with a laser pulse $E_{p}$ of duration $T_{p}$. (c) As the probe pulse travels along the fiber at light speed $c_{n}$, a point scatterer located at $z_{s}$ causes part of the incident energy $E_{i}$ to propagate back to the interrogator. (d) The backscattered wave $E_{s}$ from a single point scatterer is received around a time instant corresponding to the two-way time-of-flight between the laser source and the scatterer with a duration equal to the probe pulse. (e) An external vibration applies stress to the fiber, which deforms its shape. Under strain, the scatterers’ positions are slightly perturbed. (f) A scatterer’s displacement introduces a proportional delay to the backscattered wave. (g) The total backscattered field results from the superposition of the backscattered light from all illuminated scatterers along the fiber. (h) The time-of-flight $\tau$ of the received total backscattered signal $E_{b}$ is linearly related to the axial distance along the fiber. Cross-correlating the total backscattered signal received from adjacent fiber locations separated by the gauge length $L_{g}$, provides an estimate of the axial strain along the fiber due to external vibrations; see Sec II.5.
II.2 Rayleigh backscattered wave
DAS technology employs an optoelectronic system, referred to as the interrogator unit, to transmit a light pulse into the optical fiber and detect the power of the backscattered electric field due to the interaction of the transmitted pulse with the impurities along the fiber.Hartog2017
The most common single-mode optical fibers are designed with a cross-section of micro-metric dimensions such that only a single mode (transverse mode) of electromagnetic wave propagation is supported without modal dispersion. For a typical single-mode optical fiber, a minimum attenuation of $0.18$ dB/km is achieved for a light wavelength of $1550$ nm [Hartog2017, , Fig. 2.3], i.e., at a frequency of $200$ THz. Hence, a narrowband laser pulse at this frequency would sustain a minimum transmission loss.Miah2017 ; Hartog2017
Multimode optical fibers have a core with a larger cross-section (at least an order of magnitude larger than single-mode fibers), which guides multiple electromagnetic modes simultaneously [Hartog2017, , Sec. 2.1]. Modal dispersion of the transmitted light as it propagates along the fiber eliminates its phase coherence and polarization information. Consequently, multimode optical fibers do not support highly sensitive coherent detection methods resulting in reduced detection range and resolution. However, light emitting diodes (LED) can be used for light transmission in multimode fibers instead of highly coherent laser sources required in single-mode fibers, resulting in cost-effective systems [Grattan2000, , Ch. 1]. Despite the increased cost, coherent single-mode fiber systems are the preferred solution for DAS due to the higher sensitivity.
Figure 2 depicts the mechanism of active optic sensing in single-mode fibers and indicates its utilization for passive sensing of external vibrations. In mathematical terms, the probe pulse [Fig. 2(b)] of duration $T_{p}$, at radial frequency $\omega$ and initial phase $\phi_{0}$, generated by the laser source at $z=0$ is
$E_{p}(t)=p\left(\frac{t}{T_{p}}\right)\cos\left(\omega t+\phi_{0}\right),$ (2)
where the temporal window $p$ is defined by the apodization function (wavelet) $w$ as
$p\left(\frac{t}{T_{p}}\right)=\begin{cases}w(t),\,&\text{for}\quad\lvert t\rvert\leq\frac{T_{p}}{2}\\ 0,\,&\text{otherwise.}\end{cases}$ (3)
Consider the impulse response received at a reference location $z_{0}=0$ as a function of the travel time $\tau$ due to an ideal point scatterer at $z_{s}$ as the two-way Green’s function
$g_{s}(\tau|z_{s})=\left(g_{i}*g_{b}\right)(\tau)=\delta\left(\tau-2\frac{z_{s}}{c_{n}}\right),$ (4)
where $\delta$ denotes the Dirac delta function, * is the convolution operator and $g_{i}$ and $g_{b}$ denote the Green’s function for the incident and the backscattered wave at location $z_{s}$ reached at travel time $t_{s}$ due to an impulse source at $z_{0}=0$ and $t_{0}=0$, respectively,
$g_{i}(t_{s},z_{s}|t_{0},z_{0})\!=\!\delta\!\left(\!t_{s}\!-\!t_{0}\!-\frac{z_{s}\!-z_{0}}{c_{n}}\!\right)\!=\!\delta\!\left(t_{s}-\frac{z_{s}}{c_{n}}\right),$ (5)
$g_{b}(t_{0},z_{0}|t_{s},z_{s})\!=\!\delta\!\left(\!t_{0}-\!t_{s}\!-\!\frac{z_{0}-z_{s}}{c_{n}}\!\right)\!=\!\delta\!\left(\frac{z_{s}}{c_{n}}-t_{s}\right).$
More generally, the backscattered wave [Fig. 2(d)] from a point scatterer at $z_{s}$ with scattering strength $s(z_{s})$ due to the excitation pulse (2) is
$E_{s}(\tau|z_{s})$ $=E_{p}(\tau)*\left(g_{s}(\tau|z_{s})s(z_{s})\right)$ (6)
$=E_{p}\left(\tau-2\frac{z_{s}}{c_{n}}\right)s(z_{s}).$
External vibrations, such as an impinging acoustic wavefield at angle of incidence $\theta$ characterized by a particle displacement $\mathbf{u}$, apply stress to the fiber, which deforms its shape [Fig. 2(e)]. Under strain, any scatterer’s position is slightly perturbed by $\Delta z$, resulting in a proportional delay at the associated backscattered signal, while the scattering strength $s(z_{s})$ remains the same [Fig. 2(f)]:
$E_{s}(\tau|z_{s}+\Delta z)$ $=E_{p}(\tau)*\left(g_{s}(\tau|z_{s}+\Delta z)s(z_{s})\right)$ (7)
$=E_{p}\left(\tau-2\frac{z_{s}+\Delta z}{c_{n}}\right)s(z_{s}).$
The total backscattered field at any time instant $\tau$ results from the superposition of the backscattered light from all scatterers within the isochronous illuminated fiber length defined by the pulse length $L_{p}=T_{p}c_{n}$, which making use of (6), (2), and (3) yields
$E_{b}(\tau)$ $\!=\!\!\int\limits_{-\infty}^{\infty}\!\!E_{s}(\tau|z_{s})\mathrm{d}z_{s}=\!\!\int\limits_{-\infty}^{\infty}\!\!E_{p}\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)s(z_{s})\mathrm{d}z_{s}$ (8)
$\!=\!\!\!\!\int\limits_{-\infty}^{\infty}\!\!p\!\left(\!\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\!\right)\!\cos\!\left(\!\omega\!\left(\!\tau\!-\!2\frac{z_{s}}{c_{n}}\right)\!\!+\!\phi_{0}\!\right)\!\!s(z_{s})\mathrm{d}z_{s}.$
The received time signal of the total backscattered field is mapped to the fiber’s axial dimension as [Fig. 2(h)]:
$E_{b}(z)=\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z-z_{s})\cos\left(2k(z\!-\!z_{s})\!+\!\phi_{0}\right)s(z_{s})\mathrm{d}z_{s},$ (9)
where $k=\omega/c_{n}$ is the wavenumber and $p\left((\tau-t_{s})/T_{p}\right)=w(\tau-t_{s})$ for $\lvert\tau-t_{s}\rvert\leq T_{p}/2$ is equivalent to $p\left((z-z_{s})/L_{p}\right)=w(z-z_{s})$ for $\lvert z-z_{s}\rvert\leq L_{p}/2$ with a linear transformation of variables between time and space through the speed of light, $\tau=2z/c_{n}$. The resulting backscattered field (9) is a stochastic process resulting from the convolution of the pulse smoothing kernel with the scattering strength $s$, which is a random function of the axial dimension $z$.
II.3 Spectral analysis of the backscattered field
An alternative formulation for the backscattered field results from complex analysis. Signifying complex quantities with the accent $\widetilde{\cdot}$, the Fourier transform $\widetilde{X}(\omega)$ of a signal $x(\tau)$ is defined as [Proakis1996, , Eq. 4.1.32],
$\widetilde{X}(\omega)=\int\limits_{-\infty}^{\infty}x(\tau)e^{-j\omega\tau}\mathrm{d}{\tau},$ (10)
and the spatial Fourier transform is obtained by replacing $\omega=kc_{n}$ and $\tau=2z/c_{n}$
$\widetilde{X}(k)=\int\limits_{-\infty}^{\infty}x(z)e^{-j2kz}\mathrm{d}{z}.$ (11)
Then, the probe pulse (2) is expressed as
$\widetilde{E}_{p}(t,\omega)=p\left(\frac{t}{T_{p}}\right)e^{j(\omega t+\phi_{0})}.$ (12)
Consequently, the backscattered wave from a point scatterer (6) is
$\widetilde{E}_{s}(\tau,\omega|z_{s})$ $=\widetilde{E}_{p}\left(\tau-2\frac{z_{s}}{c_{n}},\omega\right)s(z_{s})$ (13)
$=p\left(\frac{\tau-2\frac{z_{s}}{c_{n}}}{T_{p}}\right)e^{j(\omega(\tau-2\frac{z_{s}}{c_{n}})+\phi_{0})}s(z_{s})$
and the total backscattered field due to the spatial distribution of scatterers within a pulse length (8) is
$\widetilde{E}_{b}(\tau,\omega)$ $=\int\limits_{-\infty}^{\infty}\widetilde{E}_{s}(\tau,\omega|z_{s})\mathrm{d}z_{s}$ (14)
$=\!\!\int\limits_{-\infty}^{\infty}\!\!p\bigg{(}\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\bigg{)}\!e^{j\left(\omega\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}.$
It follows that the total backscattered field expressed in the spatial dimension (9) is,
$\widetilde{E}_{b}(z,k)$ $=e^{j(2kz+\phi_{0})}\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z-z_{s})s(z_{s})e^{-j2kz_{s}}\mathrm{d}z_{s}$ (15)
$=e^{j(2kz+\phi_{0})}\left[\widetilde{W}(k)\widetilde{S}(k)\right],$
where $\widetilde{W}(k)$ and $\widetilde{S}(k)$ denote the spatial Fourier transform of the pulse window and the scattering function, respectively. For example, considering the pulse window to be a rectangular function, $\widetilde{W}(k)$ is a sinc function as detailed later in  (24). The scattering function $s(z_{s})$ is a random process with zero mean and autocorrelation function $R_{ss}$. Assuming that the scatterers in the fiber are spatially uncorrelated, $R_{ss}(z_{s})=\sigma^{2}\delta\!\left(z_{s}\right)$. Hence, the Fourier transform of the scattering function is a random process with constant spectral density as results from the Fourier transform of the autocorrelation function, $\widetilde{C}_{s}(k)=\mathcal{F}(R_{ss}(z_{s}))=\sigma^{2}$. The time-frequency (14) or, equivalently, the space-wavenumber (15) formulation expresses the total backscattered field as the short-time (spatial) Fourier transformAllen1977 of the scattering function over a window defined by the probe pulse.
II.4 Differential phase-measuring coherent optical reflectometry
DAS systems employ phase-sensitive coherent optical reflectometry methods (see Appendix) to measure the phase shift in the backscattered light from two fiber sections [Fig. 2(h)]. These differential phase measurements are a proxy for estimating strain from mechanical vibrations [Hartog2017, , Ch. 6]. There are several implementations of differential phase-measuring methods. Still, the underlying principle is mixing the backscattered signal with itself at a relative delay and detecting the phase of the resulting cross-correlation. The two signals can be mixed either in the optical domain (direct detection techniques) or in the electrical domain (coherent detection techniques) and can have the same frequency (interferometric recovery technique and homodyne detections) or slightly different frequencies (dual-pulse systems or heterodyne detection).Hartog2017 ; He2021
In any case, the resulting interferometric field is detected with coherent optical reflectometry methods (36) by mixing the backscattered light $\widetilde{E}_{b}(\tau,\omega)$ with the corresponding backscattered light $\widetilde{E}_{b}(\tau+\Delta\tau,\omega)$ at a relative delay of $\Delta\tau$. For an undisturbed fiber, the relative delay corresponds to a fixed distance along the fiber referred to as the gauge length $L_{g}$. External mechanical vibrations deform the optical fiber cable and, consequently, perturb the gauge length by $\Delta L$ resulting in an actual relative delay of $\Delta\tau=2(L_{g}+\Delta L)/c_{n}$. Hence, the interference product at the photodetector is [Hartog2017, , Eq. (6.11)]
$\widetilde{E}_{b}\left(\tau,\omega\right)\widetilde{E}^{*}_{b}\left(\tau+\Delta\tau,\omega\right)=$ (16)
$\quad\Big{(}\int\limits_{-\infty}^{\infty}\!\!p\Big{(}\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\Big{)}\!e^{j\left(\omega\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}\Big{)}\cdot$
$\Big{(}\!\!\int\limits_{-\infty}^{\infty}\!\!p\Big{(}\frac{\tau\!+\!2\frac{L_{g}+\Delta L-z_{s}}{c_{n}}}{T_{p}}\Big{)}\!e^{-j\left(\omega\left(\tau\!+\!2\frac{L_{g}+\Delta L-z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}\!\!\Big{)},$
or mapped in the space-wavenumber domain,
$\widetilde{E}_{b}\left(z,k\right)\widetilde{E}^{*}_{b}\left(z+L_{g}+\Delta L,k\right)=$ (17)
$\quad\Big{(}\!e^{j(2kz+\phi_{0})}\!\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z\!-\!z_{s})s(z_{s})e^{-j2kz_{s}}\mathrm{d}z_{s}\!\Big{)}\cdot$
$\quad\Big{(}\!e^{-j(2k(z+L_{g}+\Delta L)+\phi_{0})}\!\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!w(z\!-\!z_{s})s(z_{s})e^{j2kz_{s}}\mathrm{d}z_{s}\!\Big{)}$
$=e^{-jk(2\Delta L)}\left[\widetilde{W}(k)\widetilde{S}(k)\right]\left[\widetilde{W}(k)\widetilde{S}(k)e^{j2kL_{g}}\right]^{*}$
$=e^{-j\Delta\Phi}\widetilde{C}_{b(L_{g})}(k),$
where $\Delta\Phi=k(2\Delta L)$ and $\widetilde{C}_{b(L_{g})}(k)=\left[\widetilde{W}(k)\widetilde{S}(k)\right]\left[\widetilde{W}(k)\widetilde{S}(k)e^{j2kL_{g}}\right]^{*}$ denotes the cross-spectral density of the backscattered signal at a relative delay determined by the gauge length $L_{g}$.
The gauge length determines the amount of overlap of the backscattered signals in (16), i.e. $L_{\text{overlap}}=\max\left(L_{p}-L_{g},0\right)$. Hence, the mixed signals are correlated for a gauge length shorter than the pulse length, $L_{g}<L_{p}$, due to the autocorrelation of the pulse apodization function. The gauge length should be longer than the pulse length to measure the effect of phase modulation due to external vibrations independently of the pulse that probes the fiber.
Figure 3: Three-dimensional Cartesian coordinate system indicating the direction $\mathbf{e}$ of an impinging wavefield in terms of the azimuth angle $\phi$ and the elevation angle $\theta$. By convention, the linear optical fiber cable is aligned with the z-axis. The directional angles are detailed in the corresponding two-dimensional perspectives.
Figure 4: Two-dimensional data structure in DAS. (a) An interrogator uses a laser source to transmit a pulse train $E_{p}(t)$ into the optical fiber, and (b) receives the backscattered signal $E_{b}(t)$ due to inhomogeneities in the dielectric material. (c) Back at the laser’s position, the received time series is converted to slow time $t$ and fast time $\tau$. The time scales are determined by the acquisition sampling frequency $f_{s}$ along the fast-time axis $\tau$, and the pulse repetition frequency $f_{r}$ along the slow-time axis $t$. (d) Measurements along the fast-time correspond to the spatial dimension along the fiber, $z=\tau c_{n}/2$, and are commonly downsampled to the channel spacing $L_{h}$. Phase variations in the backscattered signal measured along distance $z$ are associated with the applied axial strain $\epsilon_{zz}$, whereas phase variations in the backscattered signal measured across time $t$ are associated with the strain rate $\partial\epsilon_{zz}/\partial t$.
II.5 Axial strain estimates
The analysis herein shows that the differential phase measurements of the light backscatter are linearly related to the axial strain of the fiber, which, in turn, is related to acoustic quantities in the environment. The subscript $a$ is introduced to indicate acoustic quantities and differentiate them from the corresponding electromagnetic quantities in the previous analysis. Namely, $\omega_{a}$ denotes the radial frequency of the acoustic wave, $k_{a}=\omega_{a}/c_{a}$ denotes the wavenumber and $c_{a}$ denotes the sound speed.
Interrogating the optical fiber cable with phase-sensitive coherent optical reflectometry methods (17) provides a distributed measurement of the differential phase $\Delta\Phi$ caused by optical path length variations $\Delta L$. Specifically,
$\Delta\Phi$ $=2k\Delta L=2\frac{2\pi(n+\Delta n)}{\lambda}L_{g}\frac{\Delta L}{L_{g}}$ (18)
$=\frac{4\pi nL_{g}}{\lambda}(1+\frac{\Delta n}{n})\frac{\Delta L}{L_{g}}=\frac{4\pi n\xi L_{g}}{\lambda}\frac{\Delta L}{L_{g}},$
where $\lambda$ is the wavelength of the probe pulse and $\xi=1+\Delta n/n$ is the Pockels coefficient that accounts for the strain-optical effect, i.e. strain applied to the fiber both changes its length and modifies its refractive index.Hartog2017 ; williams2024toward In single-mode fiberglass the Pockels coefficient is $\xi=0.79$.Lindsey2021
Considering an optical fiber cable that is coupled with an elastic medium, e.g., deployed on the seafloor, the cable’s deformation $\Delta L$ is attributed to mechanical waves impinging on it [Fig. 2(e)]. Specifically, let $\widetilde{\mathbf{u}}(t,\mathbf{r})=(A\mathbf{e})\odot e^{j(\omega_{a}t-k_{a}\mathbf{e}\odot\mathbf{r})}$ denote the particle displacement vector, defined in three-dimensional space $\mathbf{r}=[x,y,z]^{T}$, due to an acoustic plane wave with amplitude $A$. The direction of the plane wave is defined by the azimuthal angle $\phi$ and the polar angle $\theta$ measured from the normal to the cable axial direction $z$ through the directional vector $\mathbf{e}=[\cos\phi\cos\theta,\sin\phi\cos\theta,\sin\theta]^{T}$ as depicted in Fig. 3. The $\odot$ operator denotes element-wise multiplication.
The particle displacement component along the $z$ direction, $\widetilde{u}_{z}(t,z)=A\sin\theta e^{j(\omega_{a}t-k_{a}\sin\theta z)}$, is related to the axial strain as [Ewing1957, , Eq. (1-2)]
$\widetilde{\epsilon}_{zz}$ $=\frac{\partial\widetilde{u}_{z}}{\partial z}=\underset{\ell\rightarrow 0}{\lim}\frac{\widetilde{u}_{z}(t,z+\frac{\ell}{2})-\widetilde{u}_{z}(t,z-\frac{\ell}{2})}{\ell}$ (19)
$\approx\frac{\Delta L}{L_{g}}=\frac{\widetilde{u}_{z}(t,z+\frac{L_{g}}{2})-\widetilde{u}_{z}(t,z-\frac{L_{g}}{2})}{L_{g}}$
$\quad\quad\quad=\widetilde{u}_{z}(t,z)\frac{e^{-jk_{a}\sin\theta\frac{L_{g}}{2}}-e^{jk_{a}\sin\theta\frac{L_{g}}{2}}}{L_{g}}$
$\quad\quad\quad=\widetilde{u}_{z}(t,z)\frac{-j2\sin(k_{a}\sin\theta\frac{L_{g}}{2})}{L_{g}}$
$\quad\quad\quad=\widetilde{u}_{z}(t,z)\left(-jk_{a}\sin\theta\right)\frac{\sin(k_{a}\sin\theta\frac{L_{g}}{2})}{k_{a}\sin\theta\frac{L_{g}}{2}}$
$\quad\quad\quad=\!\!-jk_{a}\sin^{2}\!\theta Ae^{j(\omega_{a}t-k_{a}\!\sin\!\theta z)}\mathrm{sinc}(L_{g}k_{z}),$
where $k_{z}=k_{a}\sin\theta/(2\pi)$ and the spatial derivative is approximated by the total deformation over a finite gauge length.
Equation (18) indicates that the measured differential phase is linearly related to an estimate of the axial strain over the gauge length $\widehat{\epsilon}_{zz}=\Delta L/L_{g}$ as
$\widehat{\epsilon}_{zz}=\frac{\lambda}{4\pi n\xi L_{g}}\Delta\Phi.$ (20)
In turn, the axial strain estimate (20) is related to the particle displacement of external vibrations such as the acoustic wavefield around the optical-fiber cable through (19), which reveals two important aspects for interpreting DAS measurements.
First, the dependence of the axial strain amplitude to the square of the sine of the wavefield’s angle of incidence makes the strain measurement insensitive to broadside incidence, i.e., $\widetilde{\epsilon}_{zz}=0$ for $\theta=0$. This angular dependency is crucial in interpreting DAS data, as it influences the amplitude and detectability of acoustic signals. It also limits the effectiveness of DAS in configurations where ray paths are predominantly orthogonal to the fiber-optic cable. To overcome this limitation, helically wound cables were developed, offering broadside sensitivity and enabling the observation of signals from a wider range of angles.kuvshinov2016interaction ; hornman2017field
Second, the finite difference approximation of the spatial derivative of the displacement over the gauge length introduces directivity to the measurement defined by the term $\mathrm{sinc}(L_{g}k_{z})$. Hence, the gauge length acts as a spatial window which determines the measurement resolution; see Sec. III for details.
Some DAS systems, instead of estimating the axial strain (20) by unwrapping the differential phase measurement $\Delta\Phi$, estimate the strain rate by differentiating the phase measurement between successive pulses [Hartog2017, , Sec. 6.3.2], $\frac{\partial\epsilon_{zz}}{\partial t}$. In such systems, denoting the pulse repetition period as $T$, the axial strain rate estimate is
$\frac{\partial\widetilde{\epsilon}_{zz}}{\partial t}$ $=\frac{\partial}{\partial t}\frac{\partial\widetilde{u}_{z}}{\partial z}$ (21)
$\approx(-jk_{a}\sin\theta)\mathrm{sinc}(L_{g}k_{z})\frac{\partial\widetilde{u}_{z}}{\partial t}$
$\approx\omega_{a}k_{a}\sin^{2}\theta Ae^{j(\omega_{a}t-k_{a}\sin\theta z)}\mathrm{sinc}(L_{g}k_{z}).$
Concisely, axial strain estimates are related to the particle displacement component along the fiber, whereas axial strain rate estimates are related to the corresponding component of particle velocity.
II.6 Data acquisition
The preceding analysis indicates that setting up a DAS system requires defining a few acquisition parameters, such as the channel distance, the pulse repetition rate, the pulse width or pulse length, and the gauge length.
The optical fiber of total length $L_{f}$ is interrogated at regular discrete spatial locations determined by the sampling frequency $f_{s}$. After phase unwrapping, the measurements are downsampled to fewer spatial locations referred to as channels. The channel distance $L_{h}$ is the distance between adjacent channels.
For unambiguous measurements, a new pulse is emitted only after the previous pulse has traveled along the fiber and returned to the transmission point. Hence, the pulse repetition frequency in a DAS acquisition system depends on the length of the fiber $f_{r}=1/\tau_{\text{max}}=c_{n}/(2L_{f})$. For example, for a 100 km optical fiber cable with a typical refractive index of ca $n\approx 1.46$, $f_{r}$ cannot exceed 1 kHz.
Figure 4 illustrates the timescales of a typical DAS dataset as defined by the sampling frequency $f_{s}$ and the pulse repetition frequency $f_{r}$ [Hartog2017, , Fig. 6.1]. The output of a DAS system is structured in a two-dimensional format $E_{b}(\tau,t)$, where $\tau$ is the fast-time and $t$ is the slow-time. Measurements along the fast-time correspond to the spatial dimension along the fiber, $z=\tau c_{n}/2$, and are commonly downsampled to the channel spacing $L_{h}$. In contrast, measurements along the slow-time correspond to the temporal dimension of the dataset. Hence, the pulse repetition frequency defines the maximum observable acoustic frequency, $f_{a}^{\text{max}}<=f_{r}/2$. Note that the two-dimensional data arrangement resembles the typical data arrangement in synthetic aperture systems for radar [sar2005, , Ch. 4.6.1] or sonar applications.sas2011 In active sensing with synthetic aperture systems, though, time and space are accessed by the probe pulse and the platform motion, respectively, hence fast-time corresponds to the temporal and slow-time to the spatial dimension.
The laser pulse duration $T_{p}$, referred to as pulse width, determines the pulse length $L_{p}=T_{p}c_{n}$. The longer the pulse duration the higher the SNR, but the lower the spatial resolution as the fiber section within a pulse length acts like a single sensor. A pulse length is typically longer than the channel distance, which results in dependent channel measurements within the pulse length. Typical pulse widths are 10–1000 ns, corresponding to a pulse length $L_{p}\approx$ 2–200 m for a fiber with a typical refractive index of ca $n\approx 1.46$.
The gauge length $L_{g}$ determines the overlap between the backscattered signals originating from different locations along the fiber, which are mixed at the interrogator for differential phase measurements. Consequently, the gauge length should be larger than the pulse length for the mixed backscattered signals to be independent of the pulse coherence. Additionally, the gauge length determines the resolution of the axial strain measurement along the fiber; see Eq. (19). The greater the gauge length compared to the pulse length, the more linear the relation between differential phase and strain [Hartog2017, , Fig. 6.29]. On the other hand, the impinging wavefield should remain spatially coherent within a gauge length for meaningful measurements. To fulfill these competing requirements, the gauge length is commonly set equal to the pulse length.
Figure 5: (color online) Fourier pairs of distributed sensing parameters: Space (left) and wavenumber (right). (a) Fiber aperture with a total length of $L_{f}$, (b) spatial sampling at channel spacing $L_{h}$, (c) pulse aperture of length $L_{p}$, (d) gauge aperture of length $L_{g}$, and (e) total distributed sensing function resulting from sampling the total fiber aperture at the channel spacing (multiplication) and averaging over the pulse and gauge length (convolution).
Figure 6: (color online) DAS directivity when setting the gauge length $L_{g}$ equal to the pulse length $L_{p}$ as a function of (a) the wavenumber along the fiber axis $k_{z}$ and (b) the direction of the incident angle $\sin\theta$. (c) The directional sensitivity of the axial strain.
III Distributed vs array processing
Sensing wavefields through strain variations along the length of an optical fiber offers a distributed measurement modality, which samples densely large apertures. Large apertures are required to accurately resolve the direction of arrival of the impinging wavefield, especially at low frequencies, which signifies the importance of distributed sensing compared to apertures of a limited extent that are discretely sampled by sensor arrays.VanTrees2002
Figure 5 illustrates the aperture sampling and the directivity of the distributed sensing modality. The optical fiber with length $L_{f}$ is represented as a rectangular function $\mathit{f}(z)$ in one-dimensional space $z$ and its Fourier transform $\widetilde{F}(k_{z})$ is the sinc function in the wavenumber domain $k_{z}=k_{a}\sin\theta/(2\pi)$; Fig. 5(a1)–(a2),
$\mathit{f}(z)=\Pi(\frac{z}{L_{f}})=\begin{cases}1,\,&\text{for}\quad\lvert z\rvert\leq\frac{L_{f}}{2}\\ 0,\,&\text{otherwise},\end{cases}$ (22)
$\widetilde{F}(k_{z})=\mathcal{F}(\mathit{f}(z))=\mathrm{sinc}(L_{f}k_{z})=\frac{\sin(\pi L_{f}k_{z})}{\pi L_{f}k_{z}}.$
The differential phase of the backscattered light from a laser pulse traveling along the optical fiber is sampled at $n$ channels distributed at regular spacings $L_{h}$ along the fiber, such that $L_{f}=nL_{h}$. This sampling is represented by a Dirac comb function $h(z)$ with period $L_{h}$ and its Fourier transform is another Dirac comb function $\widetilde{H}(k_{z})$ with period $1/L_{h}$; Fig. 5(b1)–(b2),
$h(z)=\mathrm{III}_{L_{h}}(z)=\sum\limits_{i=-\infty}^{\infty}\delta(z-iL_{h}),$ (23)
$\widetilde{H}(k_{z})=\mathcal{F}(h(z))=\mathrm{III}_{1/L_{h}}(k_{z})=\sum\limits_{i=-\infty}^{\infty}\delta(k_{z}-i\frac{1}{L_{h}}),$
where $\delta$ is the Dirac delta function and the sum is over all integers $i\in\mathbb{Z}$.
As the probe pulse propagates along the fiber, it creates a moving average filter for the backscattered field over the pulse length that spans, potentially, multiple channels, $L_{p}=mL_{h}$ with $m\geq 1$. Using a pulse length at least equal to the spatial sampling distance determined by the channel spacing results in a densely populated sensing array as the ones commonly used for synthetic apertures in sonar and radar applications.Rolt1992 The pulse aperture (3) is represented here for illustration simplicity by a rectangular function $\mathit{w}(z)$ with Fourier transform $\widetilde{W}(k_{z})$; Fig. 5(c1)–(c2),
$\mathit{w}(z)=\Pi(\frac{z}{L_{p}})=\begin{cases}1,\,&\text{for}\quad\lvert z\rvert\leq\frac{L_{p}}{2}\\ 0,\,&\text{otherwise},\end{cases}$ (24)
$\widetilde{W}(k_{z})=\mathcal{F}(\mathit{w}(z))=\mathrm{sinc}(L_{p}k_{z})=\frac{\sin(\pi L_{p}k_{z})}{\pi L_{p}k_{z}}.$
Estimating strain over a finite gauge length introduces another spatial filter to the measurement of external vibrations; see (19). The gauge length can be represented as a rectangular aperture $\mathit{g}(z)$ with Fourier transform $\widetilde{G}(k_{z})$; Fig. 5(d1)–(d2),
$\mathit{g}(z)=\Pi(\frac{z}{L_{g}})=\begin{cases}1,\,&\text{for}\quad\lvert z\rvert\leq\frac{L_{g}}{2}\\ 0,\,&\text{otherwise},\end{cases}$ (25)
$\widetilde{G}(k_{z})=\mathcal{F}(\mathit{g}(z))=\mathrm{sinc}(L_{g}k_{z})=\frac{\sin(\pi L_{g}k_{z})}{\pi L_{g}k_{z}}.$
Since the differential phase measurements are unambiguously related to strain from external vibrations only for $L_{g}\geq L_{p}$ (see Sec. II.4), it is the gauge length that determines the final spatial resolution for DAS measurements.
The resulting distributed aperture $d(z)$ is constructed from sampling the optical fiber at channel spacing, $(f\times h)(z)$, and convolving the result with the pulse. Due to averaging over the gauge length, the output is convolved with (25). Combining (22)–(25) in either spatial or wavenumber domain yields
$d(z)=((f\times h)*w*g)(z),$ (26)
$\widetilde{D}(k_{z})=((\widetilde{F}*\widetilde{H})\times\widetilde{W}\times\widetilde{G})(k_{z}),$
as depicted in Fig. 5(e1)–(e2)
Setting $L_{g}=L_{p}$ results in a directivity that is determined by
$(\widetilde{W}\times\widetilde{G})(k_{z})\!=\!\mathrm{sinc}(L_{p}k_{z})\mathrm{sinc}(L_{g}k_{z})\!=\!\mathrm{sinc}^{2}(L_{g}k_{z}),$ (27)
as depicted in Fig. 6(a) or, equivalently, in Fig. 6(b) rescaling the wavenumber axis by the wavelength $\lambda_{a}$ to indicate the directional dependence $\sin\theta$. The axial strain estimate has a directional sensitivity with respect to the incident angle $\theta$, which is proportional to the square of the sine of the incident angle (19), as shown in Fig. 6(c). Therefore, the higher the frequency (smaller wavelengths) or the longer the gauge length, the more directive the DAS response to external vibrations becomes. However, the more directive the DAS response is, the less sensitive the measurement is to external fields due to the directional response of the axial strain. Low frequencies such that $\lambda_{a}\geq L_{g}$ are detected with higher sensitivity.
Figure 7: (color online) Deployment geometry of the north cable. The fiber portion corresponding to the data analyzed herein is indicated.
Table 2: Dataset parameters.
Parameter Dataset
description formula unit A, C B
refractive index $n_{f}$ 1.4682 1.4682
channel spacing $L_{h}$ m 2.042 2.042
number of channels $n_{h}$ 32600 32600
fiber length $L_{f}\!=\!L_{h}n_{h}$ km 66.57 66.57
pulse width $T_{p}$ ns 250 150
pulse length $L_{p}\!=\!T_{p}/c_{n}$ m 51.08 30.65
[…]
fast-time sampling frequency $f_{s}=c_{n}/L_{h}$ GHz 0.1 0.1
slow-time sampling frequency $f_{r}$ Hz 200 500
number of slow-time samples $n_{t}$ 12000 15000
slow-time duration $T_{rec}\!=\!n_{t}/f_{r}$ s 60 30
IV Sound source detection
The data analysis presented in this section is based on material provided by the Ocean Observatories Initiative (OOI), a major facility fully funded by the National Science Foundation under Cooperative Agreement No. 1743430, and the Woods Hole Oceanographic Institution OOI Program Office.dataOOI The analysis is based on datasets collected by the OOI Regional Cabled Array deployed offshore Central Oregon in November 2021. This region was covered in the Cascadia initiative with 60 Ocean bottom seismometers. toomey2014cascadia The data and associated metadata are publicly available by the University of Washington.
The analysis herein is based on three datasets. Datasets A: North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T022302Z and B: North-C2-HF-P1kHz-GL30m-Sp2m-FS500Hz_2021-11-03T015731Z feature acoustic signals that are attributed to whale calls and ship noise, respectively.wilcock2023distributed Marine mammal vocalizations and ship noise are loud acoustic sources that are detectable over long distances. The emanating wavefields insonify a substantial length of the fiber optic cable and can serve as sources of opportunity for inferring environmental parameters, such as sediment thickness and density.kuna2021seismic
Dataset C features lower frequency acoustic signals in the water column due to a distant earthquake. Such signals are known as T waves, i.e., Tertiary waves, as they arrive after P or S waves.okal2008generation The seismic waves from the earthquake initially propagate through the earth causing high frequencies to be filtered out and then couple into the water column on sloping bathymetry near the earthquake.degroot2001excitation The resulting T waves propagating as acoustic waves in the water column tend to be low frequency with a peak energy at frequencies below 10 Hz. To appreciate the spatio-temporal distribution of T waves, dataset C concatenates the data in four successive files, namely North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093202Z, North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093302Z, North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093402Z, and North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093502Z.
Table 2 lists the relevant acquisition parameters for all datasets. Figure 7 shows the deployment geometry of the north cable, from which the datasets above are collected, and indicates the range of the data utilized for the respective analysis.
Figure 8: (color online) (a) Recorded data in strain units (b) corresponding f-k representation indicating the dominant very low-frequency energy of ocean and seismic waves. (c) The same data after high-pass filtering featuring acoustic wavefronts propagating along the fiber and (d) the corresponding f-k representation.
Figure 9: (a) Wavenumber-time representation of the data in Fig. 8(c) and (b) the average response across time with notches at multiples of $k=1/L_{g}$.
Figure 10: (color online) Estimates of (a) the frequency $f_{a}$ (29) and (b) the angle of incidence $\theta$ (30) of the acoustic wavefield sensed by the optical fiber cable.
Figure 11: (color online) (a) The phase and (b) the spatial phase gradient of the analytical signal in Fig. 8(a) corresponding to the earliest arrival. The corresponding values for a fitted simulation example, assuming either (c), (d) plane wave propagation or (e), (f) spherical wave propagation.
Figure 12: (color online) Geometry of spherical wave propagation for an omnidirectional source at range $r$ from the point of closest approach $z_{\rm pca}$ and corresponding incidence angle as a function of axial distance $z$.
Figure 13: (color online) (a) Backpropagation of the spatial phase gradient of dataset A and (b) the corresponding spatial covariance matrix. (c) The closest point of approach is the point with the lowest spatial coherence along the diagonal of the spatial covariance matrix. (d) Location of the closest point of approach on the fiber.
IV.1 Whale calls
Dataset A was collected along the transmit fiber in the north cable on 4/11/2021. The duration of the recording is 60 s sampled at 200 Hz. The channel spacing is $L_{h}=2.0419$ m and the gauge length is set to $L_{g}=25L_{h}\approx L_{p}\approx 50$ m.
Figure 8(a) shows the time series recorded over 10000 channels (channels 10000–20000) after de-trending, i.e., subtracting the mean per recording and transforming the differential phase measurement to strain (20). Applying the two-dimensional Fourier transform to the recorded data, the frequency-wavenumber (f-k) representation in Fig. 8(b) is obtained, which shows that the recorded signal is dominated by very low-frequency ocean and seismic waveswilliams2019distributed . In Fig. 8(b), positive wavenumbers represent wave propagation from the ocean toward the shore and vice versa. Signals at the frequency range 0.2–1 Hz are attributed to seismic Rayleigh/Scholte waves with phase velocity in the range 200–1000 m/s and more energy propagating from shore toward the ocean; see Sec. VI.1. The lower-frequency signals ($<0.1$ Hz) represent ocean surface gravity and infragravity waves, with phase velocity ca. 20 m/s and more energy propagating from the ocean toward the shore; see Sec. VI.2.
To remove the dominant very low-frequency signals, the data are filtered with a digital eight-order Butterworth high-pass filterscipy with a cut-off frequency of 15 Hz. The high-pass filtered data in time-space and f-k representation, depicted in Figs. 8(c) and (d) respectively, indicate narrowband (15–20 Hz) acoustic arrivals that repeat at regular intervals (every ca. 20 s) and are classified as finback whale calls.wilcock2023distributed ; watkins1987 The f-k spectrum of the filtered data in Fig. 8(d) shows that these signals propagate in both directions along the cable at an apparent phase velocity of ca. 1500 m/s, consistent with an acoustic wave in the water column.
Figure 8(d) indicates that there are regular notches along the wavenumber dimension. To show the wavenumber pattern clearly, Fig. 9(a) shows a representation of the data in Fig. 8(c) in the time-wavenumber domain, i.e., applying the Fourier transform only along the spatial dimension, and Fig. 9(b) shows the average across time. The notches along the wavenumber dimension occur at multiples of $k=\pm 1/L_{g}\approx\pm 0.02$ m-1 in accordance to the gauge length beampattern; see Fig. 5(d2).
Applying the Hilbert transformOppenheimBook ; scipy on the real-valued two-dimensional dataset $d(z,t)$ in Fig. 8(c) provides the complex-valued analytic expression for the estimated strain (19) $\widehat{\epsilon}_{zz}(z,t)=d(z,t)*\left[\delta(t)+j/\pi t\right]$ such that the instantaneous phase is recovered, i.e.,
$\phi_{a}(z,t)=\angle{\widehat{\epsilon}_{zz}(z,t)}=\omega_{a}t-k_{a}\sin\theta z.$ (28)
The temporal gradient of the unwrapped phase (28) at each location $z_{c}$ is proportional to the acoustic signal frequency as
$\Delta_{t}\phi_{a}(z_{c},t)=\omega_{a}\Delta t=\frac{2\pi f_{a}}{f_{r}}\quad\Rightarrow$
$f_{a}=\Delta_{t}\phi_{a}(z_{c},t)f_{r}/(2\pi).$ (29)
Whereas, the spatial gradient of the unwrapped phase (28) at each time instant $t_{c}$ is proportional to the acoustic wavenumber and the direction of arrival $\theta$ defined in Fig. 3 as
$\Delta_{z}\phi_{a}(z,t_{c})=-k_{a}\sin\theta\Delta z=-\frac{2\pi f_{a}}{c_{a}}\sin\theta L_{h}\quad\Rightarrow$
$\theta=\arcsin\left(-\frac{\Delta_{z}\phi_{a}(z,t_{c})c_{a}}{2\pi f_{a}L_{h}}\right),$ (30)
where $f_{a}$ is estimated by the temporal gradient of the phase (29). Figure 10 shows (a) the acoustic frequency and (b) the incidence angle as estimated through the temporal and spatial gradient of the phase of the complex-valued data, respectively.
To highlight the characteristics of the V-shaped arrivals attributed to whale vocalizations, Fig. 11 details the phase (28) and the spatial phase gradient (30) of the earliest arrival, recorded at around 10 s, and compares the results with a fitted simulation example. The simulated wave is a sinusoidal pulse of 1 s duration, at a frequency of 18 Hz as estimated from the temporal phase gradient (29). The simulated pulse is propagating at a sound speed of 1500 m/s and the point of closest approach is set at $z_{\rm pca}=28.4$ km as inferred from the apex of the wavefront arrival in Fig. 8(c). Considering plane wave incidence at an angle of 70 as estimated from  (30), the propagation delay along the $z$ axis is calculated as $t=(z-z_{\rm pca})\sin\theta/c$. With the plane wave assumption, the simulation in Figs. 11(c) and (d) is already a good approximation of the recorded data, especially at distances far from the point of closest approach. Considering spherical, omnidirectional wave propagation from a source at a range of $r=1$ km from the point of closest point of approach the propagation delay along the $z$ axis is $t=(\sqrt{(z-z_{\rm pca})^{2}+r^{2}}-r)/c$. In this case, the simulation in Figs. 11(e) and (f) approximates the data more accurately also at distances around the point of closest approach. Figure 12 shows the wavefront angle of incidence, $\theta=\arctan\left((z-z_{\rm pca})/r\right)$, as a function of the distance from the point of closest approach $z-z_{\rm pca}$ for different source ranges $r$.
The uncertainty in the channel locations, the sound speed profile, and the low signal-to-noise ratio make source localization by backpropagation of the amplitude or phase of the recorded signal challenging. Backpropagating the spatial phase gradient $\Delta_{z}\phi$ of the recorded signal allows an accurate estimation of the location along the cable where the source has its closest point of approach (CPA). Consider a two-dimensional grid in time $t_{0}\equiv t$ and space $z_{0}\equiv z$ with potential spatiotemporal source locations $(t_{0},z_{0})$, i.e., each point on the grid represents a potential apex of a wavefront as shown in Fig. 11. Assuming plane wave propagation at $\theta=90$, a reasonable approximation for $z-z_{\rm pca}>10$ km as shown in Fig. 12, summing the data points over all channels $z$ (that is channels 10000–20000) along the wavefront delays $\left|z-z_{0}\right|/c_{a})$ populates the $(z_{0},t_{0})$ grid as,
$B(z_{0},t_{0})=\sum_{z}\Delta_{z}\phi_{a}(z_{0},t_{0}+\left|z-z_{0}\right|/c_{a}).$ (31)
The two-dimensional backpropagation result normalized to the maximum absolute value $\mathbf{B}/\max(\left|\mathbf{B}\right|)$ is shown in Fig. 13(a). Since the spatial phase gradient changes sign symmetrically around the wavefront’s normal incidence, it will sum up most destructively at the point of closest approach. Figure 13(b) shows the spatial covariance matrix $\mathbf{C}_{b}=\mathbf{B}\mathbf{B}^{T}$ of the two-dimensional backpropagation output $\mathbf{B}$ in Fig. 13(a). The point of closest approach is identified by locating the minimum of the diagonal of the spatial covariance matrix, $z_{\rm pca}=\arg\min\left(\text{diag}(\mathbf{C}_{b})\right)$; see Fig. 13(c).
Figure 14: (color online) High-pass filtered (cut-off frequency 15 Hz) data featuring a broadband signal attributed to a passing ship. (a) Recordings along successive channels in the north cable optical fiber. Two faint fin whale calls can be observed at around 10 and 20 seconds into the recording. (b) Data representation in the f-k domain. The frequency sweep, centered around 20 Hz, is attributed to the fin whale calls which overlap with the ship’s signal.
Figure 15: (color online) Estimates of (a) the frequency $f_{a}$ (29) and (b) the angle of incidence $\theta$ (30) of the acoustic wavefield sensed by the optical fiber cable.
Figure 16: (color online) (a) Backpropagation of the spatial phase gradient of dataset B and (b) the corresponding spatial covariance matrix. (c) The closest point of approach has the lowest spatial coherence along the diagonal of the spatial covariance matrix. (d) Location of the closest point of approach on the fiber.
IV.2 Ship noise
Dataset B was collected along the transmit fiber in the north cable on 3/11/2021. The duration of the recording is 30 s sampled at 500 Hz. The channel spacing is $L_{h}=2.0419$ m and the gauge length is set to $L_{g}=15L_{h}\approx L_{p}\approx 30$ m.
Similar to Fig. 8, Fig. 14 shows the time-space and f-k representation of the signal recorded over 3000 channels (channels 22500–25500) after de-trending, transforming the measurements to strain units and high-pass filtering at a cut-off frequency of 15 Hz. The recorded signal in dataset B indicates omnidirectional excitation in the frequency range 40–60 Hz, which is attributed to ship noise.wilcock2023distributed The notches along the wavenumber dimension occur at multiples of $k=\pm 1/L_{g}\approx\pm 0.033$ m-1.
Figure 15 shows (a) the acoustic frequency and (b) the incidence angle as estimated through the temporal and spatial gradient of the phase of the complex-valued data, respectively. Finally, Fig. 16 shows the result of backpropagating the spatial phase gradient $\Delta_{z}\phi$ of the recorded signal and the estimation of the closest point of approach, similar to Fig. 13.
Figure 17: (color online) T wave signal from the Mw 5.2 Fox Islands earthquake on November 4, 2021 at 08:57:06 UTC. (a) Map with great circle arrival of T-wave, (b) bathymetry and time-space figure recorded on the north cable. (c) waveform and spectrogram of a receiver at 40 km.
IV.3 Earthquake noise
T waves, i.e., acoustic waves coupled into the ocean at a sloping seafloor interface, such as a seamount or the continental margin, due to seismic body waves from nearby earthquakes,okal2008generation can be observed with underwater DAS cables at large epicentral distances.Ugalde2022 T waves from several earthquakes have been observed in the OOI DAS data.shen2024ocean Among these, the Mw 5.2 Fox Islands earthquake, which occurred on November 4, 2021, at 08:57:06 UTC, is one of the largest events recorded during this experiment.shen2024ocean
Figure 17 illustrates the T wave from this earthquake, observed 35 minutes and 54 seconds after the event on the OOI North Cable. The data are 1–30 Hz 8th order bandpass Butterworth filtered and then a bandpass f-k filter to retain only the signals with an apparent velocity between 1.4–3 km/s. An f-k filter processes seismic data by analyzing it in the frequency-wavenumber (f-k) domain. This technique allows for the isolation or suppression of specific seismic wave components based on their apparent phase velocity. To apply an f-k filter, the seismic data is first transformed from the time-space domain into the f-k domain. In the f-k domain, the seismic signal corresponding to the desired apparent velocity is retained. The apparent velocity is the frequency/wavenumber ratio, and in the f-k domain, this is a line with the slope equal to the apparent velocity. Once the filtering is applied, the data is then transformed back into the time-space domain.
The T wave is observed around 30 km along the cable, which can be due to a change of the bathymetry reducing the modal cut-off frequency. The reduction of the T wave energy at approximately 60 km along the cable is attributed to the directional sensitivity of DAS, explained in Sec. II.5, since the T wave signal arrives at a 120 heading, resulting in a horizontal incident angle of 60–70 with the cable; see Fig. 17(a).
V System design considerations
V.1 Optical multiplexing
Most DAS applications have used dark fibers, i.e., redundant network fibers not used for communication, to avoid interference, as overlapping DAS pulses with existing optical traffic can degrade both communication and DAS data quality. Most fibers in subsea networks are actively used for telecommunications, unlike terrestrial networks where dark fibers are more readily available due to lower relative installation costs. Consequently, dark subsea fibers are rare, and the high demand for intercontinental communication further limits their availability. Optical multiplexing methods based on frequency division can overcome this limitation.
Ordinary infrared optical communication typically occurs in the C-band (“Communication”), 1530–1565 nm, divided into channels. The information to be transmitted is optically encoded by a transceiver operating at one channel’s wavelength (or equivalently, frequency), and the output of all transceivers is combined into the same fiber—a technique known as wavelength domain multiplexing (WDM)—to permit simultaneous transmission of independent information. Initially, simultaneous operation of DAS and communication traffic in the same fiber was demonstrated with C-band WDM, dedicating three channels for the DAS.huang2019first ; jia2021experimental ; guerrier2022vibration The disadvantages of this approach are that the communication bandwidth is reduced relative to a fiber without DAS and the peak DAS power must be limited to prevent interference effects.jia2021experimental
Another approach utilizes WDM of C-band communications with DAS operating in the L-band (“Long”), 1565–-1625 nm, which has only slightly higher attenuation than the C-band but is used much less frequently in communications.wienecke2023new L-band multiplexed DAS has been demonstrated both in laboratory,wienecke2023new and in field experiments.brenne2024non ; shi2024multiplexed
V.2 Measurement range
Commercially available DAS systems have a measurement range up to 170 km.waagaard2021realtime Most subsea cables utilize repeaters at regular intervals every 40–100 km to amplify the optical signal and counteract transmission loss. However, common repeaters are effectively one-way gates preventing any backscattered signal from returning back to the interrogator. Consequently, the DAS measurement range is limited to the fiber length that spans the distance between the interrogator and the successive repeater.
To enable multi-span DAS, different optical architectures have been proposed, either utilizing existing repeaters with high-loss loopback (HLLB) paths that return a small fraction of transmitted light via the second fiber of a pairmazur2024real ; ronnekleiv2025range or bi-directional bespoke repeaters that amplify the backscattered light from distant spans and allow it to propagate backwards.cunzheng2022246km ; ip2022over HLLB paths were first used for interferometryMarra2022 ; mazur2022transoceanic and polarization sensing.costa2023localization Multi-span DAS on the HLLB path was demonstrated when combined with optical multiplexingronnekleiv2025range or coherent optical frequency reflectometry.mazur2024real While these technologies are promising for generalizing DAS across the global subsea cable network, currently none are commercially available and all technologies require significant customization for particular cable configurations.
V.3 High-frequency signal observation
While most studies have demonstrated the capabilities of DAS in observing hydro-acoustic signals at frequencies below 100 Hz—such as those produced by baleen whales and ship noise—the maximum frequency for DAS systems is inherently determined by the pulse repetition rate of DAS recordings, which is limited by the two-way travel time of optical pulses along the cable. For instance, a 50 km cable allows a pulse repetition frequency of $f_{r}=c_{n}/(2L_{f})$=2 kHz resulting in a maximum detectable frequency of $f_{\text{max}}=f_{r}/2=1$ kHz. However, several factors can reduce DAS sensitivity to high-frequency signals:
  1. 1.
    Signal Attenuation: Higher-frequency signals attenuate more rapidly in water and in the sediment (for buried cables), leading to weaker signals impinging on the cable. Currently, DAS is considerably less sensitive to high-frequency signals than ocean-bottom seismometers or hydrophones.
  2. 2.
    DAS directivity vs. directional sensitivity: At higher frequencies, the directivity of DAS, which is determined by the gauge length, becomes narrower and the response is severely attenuated by the directional sensitivity of linear DAS cables; see Fig. 6.
  3. 3.
    Gauge/Pulse length - SNR trade-off: Longer pulses carry more energy, which improves the signal-to-noise ratio (SNR) of the recorded signals. However, the longer the pulse width, the larger the gauge length required, since $L_{g}\geq L_{p}$ for unambiguous measurements of the external vibrational field, and consequently the more directive the DAS response resulting in reduced sensitivity to higher frequencies.; see Sec. III
Despite these challenges, several case studies have demonstrated the potential of DAS for observing acoustic signals above 100 Hz. Specifically, ship noise was observed with DAS up to 120 HzThiem2023ShipNC ; cao2023near and a sound source emitting frequencies up to 160 Hz was utilized to localize a DAS cable in shallow watershen2024high , whereas Ref. taweesintananon2021distributed, demonstrated DAS observing seismic airgun pulses up to 960 Hz. In another study,douglass2023distributed a 3.5 km subsea cable sampled at 2 kHz captured impulsive signals from a low-power acoustic source near the surface, measuring frequencies up to 700 Hz. A recent study in Lake Zurich, Switzerland, used a 1 km custom cable designed for DAS to record an acoustic source at frequencies up to 2500 Hz,rychen2023test a similar maximum frequency as demonstrated for communication signals.potter2024first
The use of DAS for measuring signals over a broader frequency range remains an active area of research, requiring a deeper understanding of the channel response and DAS sensitivities to high-frequency hydro-acoustic signals. Further advancements could significantly enhance the applicability of DAS for monitoring underwater soundscapes.
V.4 Unknown geometry
The exact channel locations in DAS are often unknown, unlike hydrophone arrays in ocean acoustics. Particularly, the precise positioning for most telecom cables is either unavailable or proprietary information. Even when the overall cable location is known, estimating the precise channel locations remains challenging. The exact speed of light in optical fiber cannot be measured in situ, leading to uncertainty in converting fast time to distance. Previous studies have demonstrated that a sound source broadcasting near the surface can localize the cable in shallow water. shen2024high By extension, ships, when combined with AIS location data, can serve as opportunistic sources to help determine the cable’s position.
VI Ocean applications beyond acoustics
Since fiber-optic cables deployed on the seafloor are subject to a wide range of perturbations, including strain, temperature, and pressure originating from ocean waves, bottom currents, earthquakes, and slope movement, this section presents applications of distributed acoustic sensing in oceanography and seismology, beyond acoustics.
VI.1 Ocean-bottom seismology
Figure 18: (color online) (a) Ambient seismic noise (Scholte waves) along the OOI RCA North cable between channels 9000–14000 (depth ca. 200 m), band-pass filtered in the frequency range 0.2–1 Hz. (b) Frequency-wavenumber spectrum of the same data segment. Positive wavenumbers indicate propagation towards the shore. (c), (d) Same as (a), (b) except for channels 27500–32500 (depth ca. 500 m) in deeper water.
The paucity of long-term seismographic stations in the oceans is a challenge in global seismology, inhibiting earthquake monitoring and early warning for offshore fault zones and limiting ray path coverage across broad regions of the core and mantle in tomographic models.kohler2020plan Fiber-optic sensing on subsea cable networks offers a promising solution to increase global coverage. DAS recordings of earthquakes on subsea cables have been utilized for event detection and location,lior2021detection ; baba2023observation identification of offshore fault zones from scattered waves,lindsey2019illuminating and quantification of linear and nonlinear site effects for strong ground motion.viens2022nonlinear Because the DAS measurement is made onshore in the interrogator, onshore-offshore telemetry occurs at the speed of light. Recent tests integrating ocean-bottom DAS with real-time earthquake early warning, suggest a significant reduction in warning time (up to 10 s) for earthquakes nucleating on offshore fault zones.romanowicz2023seafoam ; yin2023real
Ocean-bottom DAS offers a unique opportunity to probe the shear-wave velocity structure of marine sediments and oceanic crust, with good horizontal and frequency ($<$10 Hz) resolution. DAS can outperform conventional ocean-bottom seismometers as horizontal-component seismic sensors,janiszewski2023broad since the low profile of cables above the seafloor reduces the influence of ocean currents and the compliance noise from ocean waves can be removed by array processing (see below). The horizontal orientation makes DAS sensitive to shear waves from earthquakes and shear-wave resonances in seismic ambient noise, which can be used to infer the thickness of submarine sediments.trabattoni2024sediment ; taweesintananon2024near
For plane harmonic waves, the axial strain $\epsilon_{zz}$ can be expressed as the ratio between the axial particle velocity $v_{z}$ and the apparent phase velocity $c=\omega/(k_{a}\sin\theta)$
$\epsilon_{zz}=\frac{\partial u_{z}}{\partial z}=\frac{\partial u_{z}}{\partial t}\frac{\partial t}{\partial z}=\frac{v_{z}}{c},$ (32)
which is equivalent to (19).wang2018ground Hence, ocean-bottom DAS is particularly sensitive to low-phase velocity or short-wavelength phases, like Scholte waves (water-sediment interface modes). Array analysis of Scholte waves with meter-to-kilometer scale resolution permits recovery of frequency-dependent multi-modal dispersion (Fig. 18), which has been widely utilized to invert for the depth-dependent shear-wave velocity structure of shallow marine sediments. spica2020marine ; cheng2021utilizing ; williams2021scholte Ocean-bottom DAS observations of seismic noise—for example, the 0.2–1 Hz noise radiating from the coast evident in f-k spectra from the OOI RCA (Fig. 18)—has been used to characterize microseisms.williams2019distributed ; guerin2022quantifying ; xiao2022locating ; fang2022unraveling
VI.2 Ocean surface waves, nearshore processes, and tsunamis
Ocean-bottom DAS is sensitive to cable strain induced by the pressure perturbations from surface gravity waves (SGW; Fig. 19).lindsey2019illuminating ; sladen2019distributed ; williams2019distributed Recent work has demonstrated that DAS can quantify sea surface statistics, such as significant wave height and average wave period. glover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations
Water depth and gauge length determine whether SGWs can be observed by ocean-bottom DAS. All SGWs follow the dispersion relation
$\omega^{2}=gk\tanh(kh),$ (33)
where $\omega$ is angular frequency, $k$ is wavenumber, $g$ is gravitational acceleration, and $h$ is water depth (Fig. 19b). Wind waves and swell in the 0.05–0.5 Hz band therefore have wavelengths of 6–600 m in deep water ($kh\gg 2\pi$), decreasing as the waves shoal into shallow water. Choice of appropriately short gauge length and pulse width (19) is therefore important for resolving the higher-frequency components of the SGW spectrum. glover2024comparisons
The pressure field of SGWs decays exponentially with depth as a function of the wavelength, such that the ratio of the pressure at the seafloor to the pressure at the sea surface ($p_{0}=\rho gA$, where $A$ is wave height) is
$\frac{p}{p_{0}}=\frac{1}{\cosh{kh}}$ (34)
where $h$ is water depth and $k=k(\omega)$ is given by (33). In shallow water ($kh\ll 2\pi$), this hydrodynamic transfer function is approximately constant, but differs by many orders of magnitude in deeper water. Consequently, DAS can only observe wind waves and swell ($\sim 50-500$ mHz) in shallow water near the coast. Infragravity waves ($\sim 3-30$ mHz) and tsunamis ($\lesssim 20$ mHz) are observable at all depths.bromirski2002near ; taweesintananon2023distributed ; glover2024measuring ; shi2024multiplexed
Figure 19: (color online) (a) Ocean surface gravity waves observed between channels 3000–4000 on the OOI RCA North cable from dataset A (depth ca. 50 m), band-pass filtered 0.05–0.5 Hz. (b) Frequency-wavenumber spectrum of the same data segment, showing propagation towards the shore at positive wavenumbers and the theoretical SGW dispersion relation (white line).
Comparison of DAS spectra with collocated buoys and pressure sensors in four recent studiesglover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations has shown that the DAS strain measurements are linearly correlated with SGW bottom pressure over a wide variety of sea states, cable types, and installation conditions. For a cable uniformly coupled to elastic sediment, the axial strain in the cable is identical to the axial strain in the surrounding sediment,kuvshinov2016interaction ; williams2024toward suggesting that the DAS sensitivity to SGW comes from seafloor compliance, as typically seen on ocean-bottom seismometers.crawford1991seafloor For DAS on unburied cables or where the sediment is nearly fluid, sensitivity to SGW should arise from direct pressure loading on the cable through the Poisson effect, which is a function of the cable construction. glover2024measuring The sensitivity could then range between $10^{-11}$ to $10^{-8}$ strain/Pa for buried cablesdoran2016infragravity and $10^{-12}$ to $10^{-10}$ strain/Pa for unburied cables.budiansky1979pressure These expectations are generally consistent with observationally derived transfer functions, though variability in sensitivity up to 6 orders of magnitude from cable to cable remains a challenge.glover2024comparisons ; glover2024measuring ; meule2024reconstruction ; smith2023observations
DAS is inherently multi-scale, with a spatial resolution span 1–105 m. Because cables are often buried in shallow water, high-fidelity wave measurements can be made in a continuous transect from the beach to the outer shelf and are immune to strong currents or sea ice. For example, a 36-km cable across the Beaufort Shelf measured the attenuation law of SGW in sea ice and mapped its spatial variability under changing ice conditions.smith2023observations In another study, array processing was applied to separate incident and shore-reflected SGWs to assess the influence of coastal reflection on microseism generation in the Mediterranean.guerin2022quantifying DAS array measurements of SGW dispersion in shallow water have also been used to study wave-current interaction and estimate depth-averaged current speed across meter to kilometer scales.williams2019distributed ; williams2022surface ; lin2024monitoring Recent observation of infragravity waves and small tsunamisshi2024multiplexed ; tonegawa2024high ; xiao2024detection also points to a future where the dense coastal network of seafloor communication and power cables can be leveraged to increase dramatically the density of the existing sparse buoy network for real-time tsunami detection and early warning.
VI.3 Temperature sensing
DAS can directly measure changes in fiber temperature. Commonly, DAS refers to the family of technologies that measure phase changes of Rayleigh-scattered light due to variations of the optical path length due to external stimuli, regardless of whether these stimuli originate from mechanical vibrations or temperature fluctuations. The main distributed temperature sensing (DTS) is a distinct technology based on Raman scattering, see Fig. 1. The temperature sensitivity of DAS has been exploited in borehole geophysics to monitor fluid flow, where it is sometimes called distributed temperature gradient sensing.sidenko2022experimental ; titov_spe ; karrenbach2019fiber ; bradley2024
For DAS, similar to (18), the phase change over a gauge length $L_{g}$ is a function of changes in the optical path length ($\Delta L/L_{g}$) and refractive index ($\Delta n/n$), which are modified by temperature through the thermo-elastic and thermo-optic effects
$\Delta\Phi$ $=\frac{4\pi nL_{g}}{\lambda}\left(\frac{\Delta L}{L_{g}}+\frac{\Delta n}{n}\right)$
$=\frac{4\pi nL_{g}}{\lambda}\left(\alpha_{T}+\psi\right)\Delta T,$ (35)
where $\Delta T$ is a temperature change, $\alpha_{T}=\Delta L/(L_{g}\Delta T)$ is the coefficient of thermal expansion, and $\psi=\Delta n/(n\Delta T)$ is the thermo-optic coefficient, reflecting the relative change in the index of refraction with temperature.sidenko2022experimental Substituting typical values of these coefficients for silica fiber,hocker1979fiber the sensitivity to temperature is $\Delta\Phi/\Phi=7\times 10^{-6}$ 1/C, where $\Phi=4\pi nL_{g}/\lambda$. Combining this with the strain sensitivity (18), a temperature change of 1C induces the same phase change $\Delta\Phi$ as a mechanical strain of $10^{-5}$ m/m.zumberge2018measuring
Internal gravity waves and turbulence in the continuously stratified ocean cause temperature oscillations at any fixed sensor. DAS recordings from seafloor cables offshore Japanide2021very , in the Canary Islandswilliams2023fiber , and Mediterranean Seapelaez2023high have revealed low-frequency transients over time-scales from minutes to weeks, inferred to represent temperature oscillations from internal tides and turbulence. These DAS temperature observations are qualitatively consistent with observations from Raman-based DTS deployed on seafloor cables.sinnett2020distributed ; davis2020fate
For oceanographic studies, the greatest advantage of Rayleigh-based DAS over Raman-based DTS is its range, which is typically 4 to 10 times longer than commercial DTS systems, such that a DAS interrogator installed on land can monitor temperature dynamics from the shallow shelf to the abyssal ocean. Compared to DTS, DAS also has improved sensitivity to small (milli-Kelvin scale) temperature fluctuations. However, DAS can only measure relative temperature, whereas Raman-based DTS can be calibrated to measure the absolute temperature at any point along a short ($<10-30$ km) cable.sidenko2022experimental ; williams2023fiber ; sinnett2020distributed
Because DAS is a single-component sensor, temperature and strain contributions to the measured phase cannot be disentangled. In some applications, temperature transients may be distinguishable from mechanical vibrations using array analysis, as performed above for acoustic signals (e.g., Fig. 8), because internal waves and turbulence occupy characteristic space-time scales.pelaez2023high However, given the gauge lengths of DAS (1–100 m), the smallest-scale temperature oscillations associated with turbulent eddies may be resolved in time but not in space, possibly contributing to the measurement noise in seismo-acoustic DAS. Cable burial insulates DAS from temperature transients, and temperature effects can be ignored at seismic and acoustic frequencies in buried cables.williams2023fiber If bespoke cables are used, DAS can be designed to orthogonalize strain and temperature by probing multiple fibers with different thermal properties.zumberge2018measuring
VI.4 Flow-related strain
Strain measurements from fiber-optic cables that are in loose mechanical contact with the surrounding medium, such as freely suspended cables in the water column or cables lying on the seafloor, are subject to flow-induced cable vibrations, which may mask or distort acoustic signals. Ugalde2022 ; mata2023identification The flow-related vibrations are called coupling noise in borehole applications, where the DAS cable is clamped to the wall of a borehole.daley2016field In an ocean setting, such vibrations might offer a unique insight into hydrodynamics when driven by near-bottom flows.mata2023monitoring ; spingys2024distributed ; spingys2024optical
Fluid mechanics predict that flow past a cable creates vortices at either side of the cable. The pressure from each vortex drives harmonic oscillations of the cable at a frequency that is given by the Strouhal number, $\text{St}=fD/v$, where $f$ is the vibrational frequency, $D$ is the diameter of the cable, and $v$ is the speed of the flow across the cable.williamson2004vortex For a typical Strouhal number of 0.2 and a telecommunication cable diameter of 2 cm, flow speeds from 0.01–1 m/s induce vibrations at the frequency range $f=v\text{St}/D\approx$ 0.1–10 Hz on free-hanging cable segments, which can pose a challenge for seismo-acoustic studies in that frequency range.
In theory, the flow speed in the plane orthogonal to the cable can be recovered from DAS data. A pioneering studymata2023monitoring compared the vibrational frequencies of a free-hanging DAS cable segment with a collocated flow velocity measurement, demonstrating the validity of this approach. However, the dynamics of the vibrational modes are affected by the cable’s length and stiffness making flow-induced strain measurements with DAS challenging.mata2023identification Flow-induced vibrations were later exploited to infer the particle velocity of passing internal waves in a Greenland fjord.graeff2024shaken
VII Conclusions
This review provides an encompassing description of DAS technology and demonstrates its potential applications to ocean acoustics. We detail how DAS converts a fiber-optic cable into a distributed sensor of vibrational fields, such as propagating sound, substantiating that active optic sensing can be used as a proxy for passive acoustic monitoring of the environment. Based on the mathematical derivation of the underlying physics from electromagnetic to mechanical and acoustic quantities in DAS measurements, we demonstrate the effect of typical DAS acquisition parameters in signal processing. The potential of DAS technology for underwater acoustic applications, such as sound source detection, is demonstrated with the Ocean Observatories Initiative data.
Deriving the directional sensitivity of the distributed sensing modality in terms of conventional array signal processing, we explain that a limitation of current DAS systems in observing high-frequency signals stems from competing data acquisition requirements. Specifically, the maximum detectable frequency is determined by the repetition frequency of the pulse that interrogates the optical fiber, which, in turn, is inversely proportional to the measurement range. Additionally, the longer the transmitted pulse length, the higher the signal-to-noise ratio of the measurand but the less sensitive the system becomes to high-frequency vibrational fields.
The extensive coverage and robustness to harsh environmental conditions make fiber-optic sensing a considerable alternative to conventional discrete arrays of sensors. This review provides the background and insight to expand the application of DAS in monitoring the oceans.
VIII Author Declarations
The authors have no conflict of interest to disclose.
IX Data Availability
The data that support the findings of this study are publicly available by the University of Washington.dataOOI
Acknowledgements.
This work was funded by the NATO Allied Command Transformation and by the Office of Naval Research under Grant N00014-21-1-2267.
Appendix A Coherent optical reflectometry
The backscattered light at the optical fiber interrogator is detected with coherent optical reflectometry. In principle, coherent optical reflectometry mixes the backscattered signal returning from the interrogated fiber $E_{b}$ with a reference signal $E_{\text{ref}}$. The resulting signal is fed to the photodetector such that the output current $I_{pc}$ is proportional to the power of the resulting interferometric field He2021
$I_{pc}$$\propto\Re\left\{\left(E_{\text{ref}}+E_{b}\right)\left(E_{\text{ref}}+E_{b}\right)^{*}\right\}$ (36)
$=\lvert E_{\text{ref}}\rvert_{2}^{2}+\lvert E_{b}\rvert_{2}^{2}+\Re\left\{2E_{\text{ref}}E^{*}_{b}\right\}.$
The reference signal $E_{\text{ref}}$ is, usually, a replica of the transmitted pulse $E_{p}$, potentially amplified and frequency shifted by an optical local oscillator
$\widetilde{E}_{\text{ref}}(\tau,\omega+\Delta\omega)=G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\widetilde{E}_{p}(\tau,\omega),$ (37)
where $G_{\text{LO}}$ is the gain and $\Delta\omega$ the frequency shift introduced by the local oscillator, whereas $\phi(\tau)$ accounts for the time-dependent phase introduced by differences in the optical path between the reference and backscattered signal. Deriving from Eqs. (37), (12), and (14), the interference product becomes
$\widetilde{E}_{\text{ref}}(\tau,\omega\!+\!\Delta\omega)\widetilde{E}^{*}_{b}(\tau,\omega)=$ (38)
$\quad G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}p\left(\frac{\tau}{T_{p}}\right)e^{j(\omega\tau+\phi_{0})}\cdot$
$\int\limits_{-\infty}^{\infty}\!\!p\left(\frac{\tau\!-\!2\frac{z_{s}}{c_{n}}}{T_{p}}\right)\!e^{-j\left(\omega\left(\tau\!-\!2\frac{z_{s}}{c_{n}}\right)+\phi_{0}\right)}s(z_{s})\mathrm{d}z_{s}=$
$G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\!\!\!\!\!\int\limits_{-L_{p}/2}^{L_{p}/2}\!\!\!\!\!\!w(z)w(z\!-\!z_{s})s(z_{s})e^{j2kz_{s}}\mathrm{d}z_{s}=$
$G_{\text{LO}}e^{j\left(\Delta\omega\tau+\phi(\tau)\right)}\left[\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\right].$
Lowpass filtering the receiver output excludes the high-frequency signals around the nominal frequency $\omega$ and retains the signal at the difference frequency $\Delta\omega$
$I_{pc}(\tau)\propto\Re\left\{\widetilde{E}_{\text{ref}}(\tau,\omega\!+\!\Delta\omega)\widetilde{E}^{*}_{b}(\tau,\omega)\right\}=$ (39)
$\quad G_{\text{LO}}\cos\left(\Delta\omega\tau\!+\!\phi(\tau)\right)\Re\left\{\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\right\}.$
Couplers can be used to extract the phase of the detected signal [Hartog2017, , Sec. 3.3.1.1]. For example, an in-phase/quadrature (I/Q) demodulation systemwang2016coherent uses a $2\times 2$ channel coupler to output both the in-phase signal (39) and the corresponding quadrature signal by phase shifting the in-phase signal by $\pi/2$
$I_{pc}(\tau,\pi/2)\propto G_{\text{LO}}\sin\left(\Delta\omega\tau+\phi(\tau)\right)\Re\left\{\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\right\}.$ (40)
Then, the phase of the detected signal can be calculated as
$\Delta\omega\tau+\phi(\tau)=\arctan\left(\frac{I_{pc}(\tau,\pi/2)}{I_{pc}(\tau)}\right)$ (41)
$\phi(\tau)=\arctan\left(\frac{\sin\left(\Delta\omega\tau+\phi(\tau)\right)}{\cos\left(\Delta\omega\tau+\phi(\tau)\right)}\right)-\Delta\omega\tau$
Since the backscattered field is a random process the output of the photodetector varies randomly, exhibiting a speckle-like behavior. However, averaging the signal at the photodetector over several transmissions the resulting output is proportional to the power spectral density of the pulse’s apodization function $\widetilde{C}_{ww}=<\widetilde{W}(k)\widetilde{W}^{*}(k)>_{p}$
$<\!I_{pc}\!>_{p}\propto G_{\text{LO}}e^{j\Delta\omega\tau}\Re\!\left\{\!<\!\widetilde{W}(k)\widetilde{W}^{*}(k)\widetilde{S}^{*}(k)\!>_{p}\!\right\}\!.$ (42)
Narrowband pulses result in a long autocorrelation response, whereas broadband excitation, e.g. frequency-modulated pulses, results in a shorter autocorrelation function and hence improves temporal resolution or, equivalently, spatial resolution. Figure 20 illustrates that the temporal resolution of a narrowband pulse is proportional to the pulse duration (a)–(c), whereas the resolution of a broadband pulse is inversely proportional to the bandwidth (d)–(f). Long pulses are required to achieve a sufficient signal-to-noise ratio (SNR). Broadband frequency modulated pulses (chirps) achieve effective pulse compression, hence finer resolution for the same pulse duration and consequently higher SNR.
Figure 20: Transmitted laser pulse characteristics. (a) Sinusoidal pulse of duration T multiplied by a tapered cosine window, (b) its frequency spectrum, and (c) autocorrelation function. (d) Linear frequency modulated pulse of duration T and bandwidth $\Delta f$ multiplied by a tapered cosine window, (e) its frequency spectrum, and (f) autocorrelation function.
Coherent optical time-domain reflectometry (COTDR) is used with narrowband sources and increases the SNR to improve detection by introducing a substantial gain to the local oscillator. Coherent optical frequency-domain reflectometry (COFDR) is used with broadband sources producing frequency-modulated pulses to increase the SNR by increasing the pulse length without compromising the resolution.